OpenStudy (anonymous):

Points of y=f(x) are: A (-1,0) B (2,1) C (5, 0) Determine the new points for y=f(4x)

OpenStudy (anonymous):

A (-4,0) B (8,4) C (20,0)

OpenStudy (dumbcow):

Multiply domain by 4 A y = f(-4) B y = f(8) C y = f(20) there is not enough information to know the y-values

OpenStudy (anonymous):

If it helps, I found out what the original function is: -1/9(x-2)^2+1. At least I think that's right

OpenStudy (dumbcow):

yeah that would help so if you know f(x), replace x with 4x to get f(4x) --> -1/9(4x-2)^2 +1

OpenStudy (anonymous):

Apparently the y values are the same as the original, but the x values aren't what you say. I've tried (-4, 0) (8, 1) (20, 0) and I'm told its wrong.

OpenStudy (anonymous):

Apparently I'm suppose to divide by 4. Does that make sense?

OpenStudy (dumbcow):

yeah it does, im trying to work out the new equation so it should be (-1/4,0) (1/2,1) (5/4,0)

OpenStudy (anonymous):

Can you explain why it works? I don't understand.

OpenStudy (dumbcow):

yeah when you replace x with 4x in the equation as i did above -> f(4x) = -1/9(4x-2)^2 +1 well the new vertex is at point where 4x-2 = 0 because standard form is a(x-h)^2 +k 4x-2 =0 -> 4x = 2 -> x = 1/2 vertex: (1/2, 1)

OpenStudy (dumbcow):

for the others plug in 0 for y -1/9(4x-2)^2 +1 = 0 -1/9(4x-2)^2 = -1 (4x-2)^2 = 9 4x-2 = +-3 4x = +-3 +2 x = (+-3 +2)/4 x = -1/4, 5/4

OpenStudy (anonymous):

Alright, I get it now. Thanks a lot.

OpenStudy (dumbcow):

no problem